The length of the Kaufman Bridge is 2,400 feet and the length of Bruner Bridge is 3,500 feet. Which of the followin statements comparing the Kaufman and Bruner Bridges is correct? Check all that apply. The Kaufman Bridge is 31.43% shorter than the Bruner Bridge. □ The Kaufman Bridge is 1,100 feet longer than the Bruner Bridge. The Bruner Bridge is 1,100 feet shorter than the Kaufman Bridge. The Kaufman Bridge is 45.83% shorter than the Bruner Bridge. The Bruner Bridge is 31.43% longer than the Kaufman Bridge The Bruner Bridge is 45.83% longer than the Kaufman Bridge.
Answer #1

IF WE COMPARE KAUFMAN BRIDGE WITH BRUNER BRIDGE

BRUNER BRIDGE IS = (3500-2400)/2400 = 45.83% LONGER THAN KAUFMAN BRIDGE

AND IF WE COMPARE BRUNER BRIDGE WITH KAUFMAN BRIDGE

KAUFMAN BRIDGE IS = (3500 - 2400)/3500 = 31.43% SHORTER THAN BRUNER BRIDGE

CORRECT ANSWER : FIRST STATEMENT & LAST STATEMENT (THUMBS UP PLEASE)